OpenStudy (anonymous):

How do I differentiate this:

OpenStudy (anonymous):

(5) Square root ^ 3 (x)

geerky42 (geerky42):

\(5\sqrt[3]{x}\) ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Use the power rule \[ y= a x^n\\ y'= n a x^{n-1} \] In your case a=5 and n=1/3